OpenStudy (anonymous):

if x+x^-1=5, then find the value of x^4 + x^-4

OpenStudy (mokeira):

find x then substitute

OpenStudy (asnaseer):

no need for that, try expanding this and see what you get:\[\left(x+\frac{1}{x}\right)^2=\]

OpenStudy (anonymous):

is it possible to solve in an another way?

OpenStudy (anonymous):

\[x ^{2}+(1/x ^{2})\]

OpenStudy (asnaseer):

not quiet, remember that:\[(a+b)^2=a^2+2ab+b^2\]

OpenStudy (anonymous):

oh there will be an extra 2

OpenStudy (asnaseer):

yes, so you get:\[\left(x+\frac{1}{x}\right)^2=x^2+\frac{1}{x^2}+2\]therefore:\[5^2=x^2+\frac{1}{x^2}+2\]

OpenStudy (asnaseer):

now simplify to get:\[x^2+\frac{1}{x^2}=..\]then apply same rule again and you're done

OpenStudy (anonymous):

ok let me try

OpenStudy (anonymous):

ok done. thankyou very much

OpenStudy (asnaseer):

yw :)

OpenStudy (anonymous):

can we simply solve this equation ? x+1/x=5 x^2-5x+1=0 http://prntscr.com/4gafv8 @rational

OpenStudy (anonymous):

even though i liked the @asnaseer answer

OpenStudy (asnaseer):

you can, but squaring twice /seems/ simpler - I guess its a matter of taste? :)

OpenStudy (anonymous):

ik :P it was just a though, cuz we might have 2 answers ( i dint check yet :P)

OpenStudy (rational):

solving x is a good idea too as x and 1/x magically give you conjugate pairs here

OpenStudy (rational):

x^4 + 1/x^4 = 5^4 - 2

OpenStudy (anonymous):

ohh both give same solution , and one solution cool :D

OpenStudy (anonymous):

hmm n0 its 23^2-2

OpenStudy (rational):

how did u work it ?

OpenStudy (asnaseer):

if you try and solve the quadratic you get:\[x=\frac{5\pm\sqrt{21}}{2}\]both of these lead to:\[x^4+\frac{1}{x^4}=527\] If you use the method I described, then you first get:\[x^2+\frac{1}{x^2}=23\]and then:\[\left(x^2+\frac{1}{x^2}\right)^2=23^2\]therefore:\[x^4+\frac{1}{x^4}+2=23^2\]therefore:\[x^4+\frac{1}{x^4}=23^2-2=527\]

OpenStudy (rational):

exactly! solving x explicitly requires binomial expansion i guess http://www.wolframalpha.com/input/?i=1%2F2%5E4%282%285%5E4%2B%5Cbinom%7B4%7D%7B2%7D*5%5E2*21%2B21%5E2%29++%29 not sure if there is a way to avoid expansion if we solve x explicit

OpenStudy (anonymous):

solving x is quadrati equation :o